# Approximating non-Gaussian Bayesian partitions with normalising flows: statistics, inference and application to cosmology

### v2

@@ 1 Introduction
For our paper, we pursue the question of whether the differentiability of the mapping constructed by the normalising flow can be exploited, as many implementations enable auto-differentiability as a numerical method. This would be an alternative pathway to cumulants, entropies or the surprise statistic. After all, Bayes partitions are generalisations of the Bayes evidence itself, so it would be sensible to expect that similar methods as for evidence computations should be applicable. In addition, we would like to find out whether the mapping constructed by the normalising flow can be integrated into analytical calculations in an advantageous way.
@@ 2.2 Normalising flows and partition functions
Applying this to our partition functions in Equation 3 leads to a Gaussianised partition function through change of variables, effectively through an integration by parts,
@@
$g(\alpha)=\left|\det\mathrm{D}f(\alpha)\right|^{1-\frac{1}{T}}\exp\left(\frac{J_{\gamma}\theta^{\gamma}(\alpha)}{T}\right)$ (7)
@@ 2.3 Moment and entropy calculations with a flow
In the following chapters the skewness parameter $s^{\gamma}$ and kurtosis parameter $\kappa^{\gamma}$ of the posterior distribution $p(\theta^{\gamma})$ will be of interest. They are simply defined as the third and fourth standardised moments, where the standardised moments are given by
$\widetilde{\mu}_{m}=\left\langle\left(\frac{\theta^{\gamma}-\mu}{\sigma}\right)^{m}\right\rangle,$ (13)
where $\mu$ and $\sigma$ correspond to the mean and variance of the corresponding marginal distribution $\Theta^{\gamma}$. The Helmholtz free energy can also be determined by the partition function and is given by
@@ 2.4 Flow expansion through a differentiation operator
Here, $g(\alpha)$ is defined as in Equation 7, and we will refer to this result as the flow expansion.
@@ 3 Application to supernova data
with $H(a|\Omega_{m},w_{0})$ denoting the Hubble function $H(a,\Omega_{m},w_{0})$ which is of the following expression for a constant equation of state parameter $w_{0}$:
@@
with $N$ being the number of samples. This requires a distributional form of the posterior $p(\theta|y)$, which is not analytically known. For comparison, we estimate $p(\theta^{i}|y)$ on a grid (effectively forming histogram) and by kernel density estimation (KDE), using the SciPy package. Those approximations are not needed when working with the normalising flow: It provides a function $f(\theta)$ that maps the samples to a standard normal distribution $p(\alpha)$, with known analytical form. Thus, by sampling from a normal distribution and using the trained normalising flow, the entropy can be computed via
@@
Table 2: Comparison of moments derived from sampling and via the flow expansion for the supernova data. Additionally, the sample estimates on the emcee posterior samples are given. Not only agree the values nicely, but also the results match the usually obtained ones (e.g. Suzuki et al., 2012).
This shows that the via the flow expansion statistical analysis of real world data is possible and agrees within its errors with our expectation. The present limitation comes from the fact that higher derivatives of normalising flows are not possible yet and smoothness can be improved. Nevertheless, also more basic methods as information entropy calculation become easier as one is able to insert known analytical distributions and just uses the Jacobian of the transformation performed by the neural network.
4 Geometry of the Normalsing Flow
@@
Calculating the coefficients introduced in Equation 25 for this matrix and rounding up to four digits, yields $\kappa=-0.0509$, $\gamma_{1}=0.0111$, $\gamma_{2}=-0.0564$, and $\omega=-0.0780$. Additionally, we note that the flow conserves orientation of the coordinate frames. The areas of cells in parameters spaces in $\theta$ and $\alpha$ are naturally related by $\kappa^{2}$, as verified in Table 3.
@@ 5 State variables $T$ and $J_{\gamma}$ of Bayesian partitions
For example, Figure 4 shows the logarithm of the partition function as a function of temperature, similarly to Kuntz et al. (2024), where a KDE approach (displayed in orange) was used. One can nicely see that both approaches agree almost perfectly. As expected, the value of $\ln Z[T]$ saturates for high temperatures $T$.
Figure 4: Plot of the partition function $\ln Z$ as a function of temperature $T$ comparing the flow (blue) result to KDE (orange) estimate as in Kuntz et al. (2024). The two results agree perfectly.
@@ Acknowledgements
Our Python implementation of the code computing cumulants and entropies from the normalising flow is available on request.
@@ Appendix A Toy model
The flow expansion with a series of derivatives defined in subsection 2.4 is applied to the toy model to calculate in each dimension the mean, variance, skewness as well as kurtosis and the covariance of both dimensions, i.e. the off-diagonal entry in the covariance matrix. The results are shown in Table 7: First, the first dimension is not only closer to the correct value, but also its error is smaller than obtained for the second dimension. This is related to the procedure how the normalising flow is constructed within FrEIA. There, the second dimension uses the output from the first dimension (Ardizzone et al., 2018) which leads to a propagation of error. One could avoid that problem by training the normalising flow a second time and swapping the input dimensions. Additionally, it is apparent that the network has learned a transformation of a Gaussian distribution to another Gaussian distribution as the values for skewness and kurtosis are especially in the first dimension very precise and exactly what is expected for a Gaussian distribution - independently of its mean and covariance: The normalising flow has to be able to reproduce a principal value decomposition which is a mere linear transform between the random variables. The obtained value for the covariance of $\theta_{1}$ and $\theta_{2}$ still agrees within its error with the ground truth, but the larger error and deviation originates from the usage of the second dimension in the calculation.
@@
Throughout this paper, we used the flow expansion up to fourth order, i.e. terminating the series defined in Equation 18 at $k=4$. This is due to the fact that the higher derivatives of the network become more and more difficult, leading to divergences. This is deeply connected to software architecture. Even for the compulsory choice of smooth activation functions, other aspects of the algorithm are not smooth and thus do not allow for a high number of derivatives. For example, the performed permutation in the AllInOneBlock is found to be very important for training, but impacts derivatives negatively. Normalising flows and as an example FrEIA are capable of learning far more complicated distributions, but for a higher complexity more layers are need. From our experience, this results in issues with higher-order derivatives of the normalising flow. But if given a smooth normalising flow which is sufficiently often differentiable, one would be able to apply the flow expansion to any order and to an arbitrary complex distribution. There are first attempts to create smooth normalising flows like for example Köhler et al. (2021), and it would be interesting to evaluate their performance beyond second order.

### v3

@@ 1 Introduction
For our paper, we pursue the question of whether the differentiability of the mapping constructed by the normalising flow can be exploited, as many implementations enable auto-differentiability as a numerical method. This would be an alternative pathway to cumulants, entropies or the surprise statistic. After all, Bayes’ partitions are generalisations of Bayes’ evidence itself, so it would be sensible to expect that methods similar to those used for evidence computations should be applicable. In addition, we would like to find out whether the mapping constructed by the normalising flow can be integrated into analytical calculations in an advantageous way.
@@ 2.2 Normalising flows and partition functions
Applying this to our partition function in Equation 3 leads to a Gaussianised partition function through change of variables, effectively through an integration by parts,
@@
$g(\alpha)=\left|\det\mathrm{D}f(\alpha)\right|^{1-\frac{1}{T}}\exp\left(\frac{J_{\gamma}\theta^{\gamma}(\alpha)}{T}\right)\,.$ (7)
@@ 2.3 Moment and entropy calculations with a flow
In the following chapters the skewness parameter $s^{\gamma}$ and kurtosis parameter $\kappa^{\gamma}$ of the posterior distribution $p(\theta^{\gamma}|y)$ will be of interest. They are simply defined as the third and fourth standardised moments, where the standardised moments are given by
$\widetilde{\mu}_{m}=\left\langle\left(\frac{\theta^{\gamma}-\mu}{\sigma}\right)^{m}\right\rangle\,.$ (13)
Here, $\mu$ and $\sigma$ correspond to the mean and variance of the corresponding marginal distribution $\Theta^{\gamma}$. The Helmholtz free energy can also be determined by the partition function and is given by
@@ 2.4 Flow expansion through a differentiation operator
Here, $g(\alpha)$ is defined as in Equation 7, and we will refer to this result as the flow expansion. At this point, we want to emphasise that solving the integral facilitates the analytical form and thus makes further statistical and respectively thermodynamic inspired calculations easier. In Sect. 3, we will comment on its numerical performance.
@@ 3 Application to supernova data
with $H(a|\Omega_{m},w_{0})$ denoting the Hubble function $H(a,\Omega_{m},w_{0})$ which is of the following expression for a constant, in terms of the scale factor $a$, equation of state parameter $w_{0}$:
@@
with $N$ being the number of samples. This requires a distributional form of the posterior $p(\theta|y)$, which is not analytically known. For comparison, we estimate $p(\theta^{i}|y)$ on a grid (effectively forming a histogram) and by kernel density estimation (KDE), using the SciPy package. Those approximations are not needed when working with the normalising flow: It provides a function $f(\theta)$ that maps the samples to a standard normal distribution $p(\alpha)$, with known analytical form. Thus, by sampling from a normal distribution and using the trained normalising flow, the entropy can be computed via
@@
Table 2: Comparison of moments derived from sampling as described in Equation 11 and via the flow expansion (Equation 18) for the supernova data. Additionally, the sample estimates on the emcee posterior samples are given. Not only agree the values nicely, but also the results match the usually obtained ones (e.g. Suzuki et al., 2012).
This shows that the via the flow expansion statistical analysis of real world data is possible and agrees within its errors with our expectation. Concerning numerical performance, the accuracy of the flow expansion for the mean and variance is a little worse than the sampling method, but still in the same order of magnitude. Both methods are roughly equally fast. The goal of the flow expansion is not to be numerically more performant, but to offer analytical improvements for statistical partition functions. The present limitation comes from the fact that higher derivatives of normalising flows are not possible yet and smoothness can be improved. Nevertheless, also more basic methods as information entropy calculation become easier as one is able to insert known analytical distributions and just uses the Jacobian of the transformation performed by the neural network.
4 Geometry of the Normalising Flow
@@
Calculating the coefficients introduced in Equation 25 for this matrix and rounding up to four digits, yields $\kappa=-0.0509$, $\gamma_{1}=0.0111$, $\gamma_{2}=-0.0564$, and $\omega=-0.0780$. Additionally, we note that the flow conserves orientation of the coordinate frames. The areas of cells in parameter spaces in $\theta$ and $\alpha$ are naturally related by $\kappa^{2}$, as verified in Table 3.
@@ 5 State variables $T$ and $J_{\gamma}$ of Bayesian partitions
For example, Figure 4 shows the logarithm of the partition function as a function of temperature, similarly to Kuntz et al. (2024), where a KDE approach (displayed in orange) was used. For simplicity, $J$ is set to zero. One can nicely see that both approaches agree almost perfectly. As expected, the value of $\ln Z[T]$ saturates for high temperatures $T$.
Figure 4: Plot of the partition function $\ln Z$ as a function of temperature $T$ with $J=0$ comparing the flow (blue) result to KDE (orange) estimate as in Kuntz et al. (2024). The two results agree perfectly.
@@ Acknowledgements
Our Python implementation of the code computing cumulants and entropies from the normalising flow is available on GitHub.
@@ Appendix A Toy model
The flow expansion with a series of derivatives defined in subsection 2.4 is applied to the toy model to calculate in each dimension the mean, variance, skewness as well as kurtosis and the covariance of both dimensions. The results are shown in Table 7: First, the first dimension is not only closer to the correct value, but also its error is smaller than obtained for the second dimension. This is related to the procedure how the normalising flow is constructed within FrEIA. There, the second dimension uses the output from the first dimension (Ardizzone et al., 2018) which leads to a propagation of error. One could avoid that problem by training the normalising flow a second time and swapping the input dimensions. Additionally, it is apparent that the network has learned a transformation of a Gaussian distribution to another Gaussian distribution as the values for skewness and kurtosis are especially in the first dimension very precise and exactly what is expected for a Gaussian distribution - independently of its mean and covariance: The normalising flow has to be able to reproduce a principal value decomposition which is a mere linear transform between the random variables. The obtained value for the covariance of $\theta_{1}$ and $\theta_{2}$ still agrees within its error with the ground truth, but the larger error and deviation originates from the usage of the second dimension in the calculation.
Throughout this paper, we used the flow expansion up to fourth order, i.e. terminating the series defined in Equation 18 at $k=4$. This is due to the fact that the higher derivatives of the network become more and more difficult, leading to divergences. This is deeply connected to the software architecture. Even for the compulsory choice of smooth activation functions, other aspects of the algorithm are not smooth and thus do not allow for a high number of derivatives. For example, the performed permutation in the AllInOneBlock is found to be very important for training, but impacts derivatives negatively. Normalising flows and as an example FrEIA are capable of learning far more complicated distributions, but for a higher complexity more layers are need. From our experience, this results in issues with higher-order derivatives of the normalising flow. But if given a smooth normalising flow which is sufficiently often differentiable, one would be able to apply the flow expansion to any order and to an arbitrary complex distribution. There are first attempts to create smooth normalising flows like for example Köhler et al. (2021), and it would be interesting to evaluate their performance beyond second order.